# A dark matter hail: Detecting macroscopic dark matter with asteroids, planetary rings, and craters

### v1

@@ I Introduction
[…]
[lines=5,findent=3pt]Macroscopic, composite, or ultraheavy objects are some of the oldest and most intriguing dark matter (DM) candidates. Early macroscopic dark matter candidates included primordial black holes [1, 2, 3, 4, 5, 6] (PBHs) and Standard Model quark nuggets [7, 8, 9], composite objects composed of up, down and strange quarks in a stable configuration and naturally formed in a strongly first order quantum chromodynamics (QCD) phase transition. While one must turn to more worked scenarios to motivate quark nuggets today [10, 11, 12], many more generic macroscopic dark matter scenarios have been studied on similar principles. For example, a global U$(1)$ symmetry can stabilize large solitons consisting of fermions or scalars, known as Fermi balls [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25] and Q-balls [26, 27, 28, 29, 30, 31] respectively.
@@
Figure 1: Constraints on dark matter from this work. In indigo, Kuiper belt objects above $\simeq 40$km in radius are evaporated from their orbits around the sun by one or more collisions with the dark matter. In blue, S-type asteroids are catastrophically destroyed by collisions with the dark matter. The dashed curve below shows the possible improvement in constraints if we observed small mass asteroids much older than their collisional lifetimes. In yellow, ice particles in Saturn’s rings are destroyed by collisions, and in salmon, planetary ring particles are evaporated from their orbits around Saturn. The olive-green region with dashed border indicates where the dark matter creates more craters on Earth than expected from near-Earth objects, but cannot be confidently considered a constraint. For reference, the densities of ice, nuclear matter, and black holes are indicated on the plot. Existing constraints on dark matter are shown in grey, as collected in Ref. [32] and cited in Section VI. On the bottom, ticks indicate how often dark matter of that size would impact the Earth. The largest previously-unconstrained region is on the larger mass end of this plot, spanning several orders of magnitude both horizontally and vertically. However, there is also a somewhat hard-to-see region above the ancient mica constraints which was also previously unconstrained [33].
@@ II Energy transfer
First we need to estimate the effect of the collision between a dark matter bullet of mass $m$ and radius $r$ and a regular object with mass $M$ and radius $R$. We will first assume that the dark matter is not destroyed or otherwise majorly affected by the impact. When the dark matter collides with the object, it will elastically scatter with particles in a tube of radius $r$ as it passes through ,until it is either stopped in the material or exits out the other side. As a simplified way to average over impact angles, we will take the tube length to be the average secant length in a sphere, $\frac{4}{3}R$, so that the volume of this tube is $V=\frac{4}{3}\pi Rr^{2}$. This also sets a limit on the maximum dark matter radius $r$ such that the dark matter still passes cleanly through the object:
@@
Enforcing this ensures the geometry of the collision is sufficiently simple, and for the scenarios we consider here $r>0.25R$ occurs in regions of parameter space which are already ruled out by other dark matter constraints and can thus be safely ignored. In principal there is no reason that we cannot look at $r>0.25R$ to extend our limits upwards, but more work would be required for the unusual geometry here.
The number of potential collisions $N$ the dark matter encounters is given by $N\simeq\rho_{o}V/m_{p}$, where $\rho_{o}$ is the density of the object and $m_{p}$ is the mass of a proton. For simplicity let us assume that the particles in the object are stationary with respect to the dark matter and treat each collision as a linear elastic collision. The change in speed after one such collision is simply,
@@
We then say that if $X>1$, the dark matter can be treated as having been stopped within the object. This gives us a lower bound on dark matter radius $r_{\mathrm{capt}}$ which can be said to be ‘captured’ by the object:
$r_{\mathrm{capt}}>$ $0.003~{}cm\times$
×$\times\left(\frac{3}{4\pi}\right)^{1/2}\left(\frac{m}{g}\right)^{1/2}\left(\frac{R}{km}\right)^{-1/2}\left(\frac{\rho_{o}}{gcm^{-3}}\right)^{-1/2}~{}.$ (4)
The energy $E$ imparted to the object by this collision can then be estimated by taking the difference between the kinetic energy of the dark matter before and after the impact. If the dark matter is captured, this is simply the total kinetic energy before the impact. If the dark matter is not captured, we then estimate,
@@
We will say that when the dark matter is at a distance $x$ throughout the tube, any particles found within the length $y$ from the center of the dark matter sphere will be ejected into space at the initial dark matter velocity. This will only serve as an extremely rough estimate for the complicated physics of the real shock which would precede the dark matter. We assume that any material farther than $y$ will deposit their energy into the material of the object, contributing to its destruction. With some geometrical working, the volume of particles lost can be shown to be,
$\int_{0}^{4/3R}\mathrm{d}x\pi y^{2}=\pi r^{3}\tan^{-1}(R/r)~{},$ (6)
and so the fraction of energy lost by this ejecta is simply,
@@ III Catastrophic Destruction
The study of the catastrophic destruction of asteroids has a long and rich history [34, 35, 36, 37, 38, 39, 40]. This is not only motivated by a desire to better understand how to avoid Earth-impacting events [41, 42, 43], but also to understand the history and formation of our solar system [44, 45, 46, 47, 48, 49]. The generic criteria for asteroid destruction is quantified with the ratio $Q^{*}$, the kinetic energy of the impacter divided by the target mass. For catastrophic asteroid destruction, $Q^{*}_{d}$ is defined as the threshold above two conditions hold. Firstly, the largest fragment should be smaller than half the mass of the original, and secondly, less than half of these fragments should eventually re-accumulate. We will adopt this criteria for destruction.
There are two important regimes for asteroids. Below roughly $0.5-1$km, asteroids are considered to be in the material strength regime, where the primary forces which must be overcome to destroy them relate to the material strength of the rock or ice. Above this, the asteroids are gravity dominated, and the energy required to overcome the gravitational binding energy is significantly larger than the material strengths.
However, destruction of gravity dominated asteroids is generally more difficult than merely overcoming the gravitational binding energy. Larger asteroids (from about $200$m to $10$km) are now thought to be predominantly rubble-pile asteroids [50], gravitational bound collections of rocks and boulders of varying sizes and compositions. Evidence for the ubiquity of rubble-pile asteroids are based on observations of low asteroid spins above about one kilometer in radius [51, 52], direct measurements of densities and compositions of individual asteroids [53, 54, 55], the observed breakup of Shoemaker-Levy 9 over Jupiter [56, 57, 58], and the DART mission [59, 60, 61, 62, 63] which intentionally collided with the asteroid Dimorphos. Since rubble pile asteroids can be highly porous, the energy of the impacter is distributed inefficiently, resulting in larger $Q_{d}^{*}$ than a naive estimate which only requires the gravitational binding energy be overcome. In general, impacts on rubble-pile asteroids must be simulated numerically, and there is a high degree of variance in outcomes depending on the propagation of shocks through the specific geometry and composition of the rubble pile.
@@
S-type $10^{3}$ $10^{6}$ $10^{-3}$ $500$ $-0.33$ $1.65$
C-type $2\times 10^{3}$ $4\times 10^{5}$ $10^{-3}$ $500$ $-0.25$ $1.23$
@@
Although the density of asteroids is both size-dependent and has some variance amongst individuals, we used the benchmark values from Ref. [64]. For S-type asteroids we took the density to be $3.0$ g cm-3, and for C-types we used $1.7$ g cm-3. For the particles in Saturn’s rings we used the density of ice $0.92$ g cm-3, but with the best-fit $Q_{d}^{*}$ for S-type targets—while these are of course not designed for a pure ice target, earlier studies [35] show that the destruction threshold for ice and basalt do not differ by more than a factor of approximately two, with ice being the easier of the two to destroy.
The smallest observed asteroids are about $2$m large, such as TC25 [65]—although this is not strictly an S- or C-type asteroid, we will use it as the benchmark smallest size. The largest S-type asteroids are Eunomia and Juno [64], around $270$ and $250$km respectively, whereas the largest C-type are Hygiea at around $430$km and Ceres at $940$km [64]. However, our constraints are not very sensitive to the largest mass objects, since the destruction criteria ceases to be met before we reach such massive asteroids.
The composition of Saturn’s rings [66, 67, 68] vary by subdivision, but generally the smallest observed particles are around a few mm and the largest on the order of $10$m. To the best of our knowledge, there is no current evidence for smaller particles than these in stable orbits [69, 70], but if there was, the constraints here could be improved. We use $1$mm as the small-size cutoff for Saturn’s ring particles, based on the treatment of monomers in Ref. [71] which we will discuss in more detail in the following subsection.
Since $Q_{d}^{*}$ is generally smaller for S-type asteroids, they tend to provide marginally tighter constraints on the dark matter than for C-types. We thus ultimately use the S-type asteroid formula when deriving the constraints. Eq. 8 holds until the objects are smaller than about $1$ m, below which the curve takes a constant value. In Ref. [34] a cutoff on this estimate is imposed at $1$cm, but in order to consider the smallest particles in Saturn’s rings, we will naively extend this to $1$mm.
@@
$t^{-1}\simeq n_{DM}\sigma v_{DM}^{2}~{},$ (11)
where we take the collision cross section to be $\sigma=\pi\left(R+r\right)^{2}\simeq\pi R^{2}$. For a given object of mass $M$, radius $R$, and lifetime $\tau$, if $t<\tau$ and the energy of the impact is sufficient to destroy it, we consider the dark matter constrained. These constraints are plotted on Fig 1.
@@ III.1 Ages of asteroids and ring particles
The observed ages of asteroids and planetary rings must then be carefully considered in order to place constraints. For larger asteroids, observations of color [72] consistently confirm ages in the billions of years, and indeed rubble-pile asteroids are very difficult to destroy by collisions with other asteroids [63]. For smaller asteroids, to the best of our knowledge, data on age is scarcer. We then must turn to simulations of the evolution of the asteroid belt from Ref. [47]. The smaller the asteroid, the easier they are to catastrophically destroy by collisions with other asteroids. By simulating the entire population, it is then possible to determine the average collisional lifetime of an asteroid of specific size, as given in Table 2. We interpolate this data to determine the lifetime $\tau$ which should be compared against the destruction timescale $t$.
@@
Table 2: Relation between radius of asteroid and collision lifetime from Ref. [47].
Of course, this lifetime is just the average lifetime of an asteroid of a given size. It is presumably the case that some of the small asteroids observed today are significantly older than the ages given here, which would substantially improve the ultimate dark matter constraints. To show this, the final constraint plot includes the region that would be constrained if we did indeed find small asteroids which could be proven to have survived $1$ Gyr without major disruptions. This is plotted as a dashed boundary in Fig. 1.
@@ III.2 Conservativeness of estimate
$E_{\mathrm{vap.}}/E_{k}=\frac{2H}{\rho_{\mathrm{granite}}v^{2}}\simeq 0.6~{},$ (12)
where $\rho_{\mathrm{basalt}}=2.8$g cm-3 and the standard bullet test speed is $v=5.5$km/s. The energy lost to vaporization (if completely vaporized) can therefore be a sizable fraction of the total kinetic energy of the impacter. In our scenario, in contrast, no energy would be lost to this vaporization.
Estimating the energy losses due to the geometry of the shocks is far more complicated, but we can get at least a rough sense of the estimation. One historically useful criteria [34, 35] for asteroid destruction used the fact that the velocity of the material shocks dropped with the distance squared in the material as the shock spread from the impact point. If the bullet impacted on one side of the material, one could then calculate how fast the shock is moving on the opposite side of the asteroid, at a distance $2R$, and if the velocity at that distant point was still larger than the escape velocity then the asteroid would be considered disrupted. In our case, the energy is not just deposited at the impact site, but throughout the secant the dark matter bullet carves out. The most distant point from the average secant we considered here is perpendicular from its midpoint, at a distance approximately $1.75R$. Then the velocity of the shock at this point will be $1.3$ times the velocity of the shock at $2R$. If we instead considered a trajectory through the middle of the asteroid (rather than the average secant), the velocity would be four times larger. Combined with the previous estimate of the energy lost to vaporization, then, we could be underestimating the constraints by an order of magnitude or so.
@@ IV Orbit evaporation
We want to know how many collisions $k$ are required to increase the velocity by a factor $\delta$. For evaporation, escape velocity at a given radius is only a factor $\delta=\sqrt{2}$ larger than the speed of circular orbit, so $\delta=\sqrt{2}$ will be the key threshold. We will use this threshold as a conservative way to place constraints, since the evaporation of objects from their orbits would trivially contradict observations. However, it is probably the case that we could place much more sensitive constraints than this—even a small disruption to the trajectories of Saturn’s rings, for instance, should have large observable consequences. We will leave this more sensitive computation to future work.
We will ignore for the moment the dark matter wind, so that the directions of the collisions are isotropic, and assume the dark matter has a uniform velocity with magnitude $v_{DM}$. The number of collisions over some timespan $\tau$ is simply,
$k=\frac{\rho_{DM}}{m}\sigma v_{DM}\tau~{},$ (13)
where we again take the cross section $\sigma=\pi\left(R+r\right)^{2}\simeq\pi R^{2}$.
@@
$\vec{v}_{f}=\vec{v}_{o}+v_{DM}\frac{m}{M}\left(1-e^{-X}\right)\sum_{k}\hat{r}_{i}~{},$ (14)
Where $\hat{r}_{i}$ are unit vectors in random directions and $\vec{v}_{o}$ is the object’s velocity. We will then make use of the handy result for the expected value of the magnitude of the vector sum of $k$ random, isotropic unit vectors:
$\langle\Big{|}\Big{|}\sum_{k}\hat{r}_{i}\Big{|}\Big{|}\rangle=\sqrt{k}~{}.$ (15)
Before proceeding, it is interesting to point out the effect of this square root on the calculation. One might naively expect that this estimate should be independent of dark matter mass, since a smaller mass implies less momentum transfer but proportionately more collisions. However, averaging over the isotropy of the dark matter and the resulting random walk of the object’s momentum is responsible for this $\sqrt{k}$ growth which ultimately breaks the degeneracy in dark matter mass.
In order to safely make the assumption of isotropy, the timescales in question should be at least on the order of $250$ million years, which is approximately the time it takes for the Solar System to traverse the Milky Way once. This means that the dark matter wind is averaged out and there is no preference for the direction of collisions. The dark matter wind can have a profound impact, since it increases both the relative velocity of collisions in the forward direction as well as the rate of such collisions—momentum transfer is therefore substantially dominated by collisions in the direction of the wind. This is particularly relevant to note for the case of Saturn’s rings, which have only been around for $100$ Myr and so have only seen a shift in the angle of the dark matter wind of somewhat less than $180^{\circ}$. For the purposes of this estimation we will ignore this effect—this will then comprise a conservative limit, since accounting for the effect of the dark matter wind would allow the ring particles to be evaporated more quickly.
@@
$\sqrt{k}$ $\geq\frac{M}{m}\frac{v_{o}}{v_{DM}}\frac{\delta-1}{\left(1-e^{-X}\right)}~{}.$ (16)
$\simeq\frac{v_{o}}{v_{DM}}\frac{R^{2}}{r^{2}}\frac{\delta-1}{2}~{}.$ (17)
Where the second line takes $X<1$. Using the relation for $k$ allows us to estimate a lower limit on the dark matter radius such that the orbit will be eventually evaporated:
$r^{2}\geq R\left(\frac{m}{\pi\rho_{DM}v_{DM}\tau}\right)^{1/2}\frac{v_{o}}{v_{DM}}\frac{\delta-1}{2}~{}.$ (18)
Now let us return to the regime when $X>1$. In this limit the dark matter is captured in the object and all of the momentum is transferred. For our estimation we will consider this a perfectly inelastic collision, with no losses from cratering or ejecta. The calculation follows similarly to the previous section, with the final velocity now being given by,
$\vec{v}_{f}=\frac{1}{M+km}\left(M\vec{v}_{o}+mv_{DM}\sum_{k}\hat{r}_{i}~{}\right),$ (19)
@@
$\frac{\sqrt{k}\frac{m}{M}}{\delta\left(1+k\frac{m}{M}\right)-1}\geq\frac{v_{o}}{v_{DM}}~{},$ (20)
which does not have a beautiful solution for $k$. We can proceed by noting that for the objects we have in question, $km<M$ is very easy to satisfy, since this limit can be simplified to,
$R>3/6\times 10^{-6}~{}\mathrm{cm}\left(\frac{\rho_{o}}{1\mathrm{gcm}^{-3}}\right)^{-1}\left(\frac{\tau}{1\mathrm{Gyr}}\right)~{},$ (21)

### v2

@@ I Introduction
Macroscopic, composite, and ultraheavy objects are some of the oldest and most intriguing dark matter (DM) candidates. Early macroscopic dark matter candidates included primordial black holes [1, 2, 3, 4, 5, 6] (PBHs) and Standard Model quark nuggets [7, 8, 9], composite objects composed of up, down and strange quarks in a stable configuration and naturally formed in a strongly first order quantum chromodynamics phase transition. While one must turn to more worked scenarios to motivate quark nuggets today [10, 11, 12], many more generic macroscopic dark matter scenarios have been studied on similar principles. In particular, a global U(1)(1) symmetry can stabilize large solitons consisting of fermions or scalars, known as Fermi balls [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25] and Q-balls [26, 27, 28, 29, 30, 31] respectively.
@@
Figure 1: Constraints on the spin-independent dark matter–nucleon cross section as a function of dark matter mass. In indigo, Kuiper belt objects above $\simeq 40$km in radius are evaporated from their orbits around the sun by one or more collisions with the dark matter. In blue, S-type asteroids are catastrophically destroyed by collisions with the dark matter. The dashed curve below shows the possible improvement in constraints if we observed small mass asteroids much older than their collisional lifetimes. In yellow, ice particles in Saturn’s rings are destroyed by collisions, and in salmon, planetary ring particles are evaporated from their orbits around Saturn. The olive–green region with dashed border indicates where the dark matter creates more craters on Earth than expected from near-Earth objects, but cannot be confidently considered a constraint. For reference, the densities of ice, nuclear matter, and black holes are indicated on the plot. Existing constraints on dark matter are shown in grey, as collected in Ref. [32] and cited in Section VI. On the bottom, ticks indicate how often dark matter of that size would impact the Earth. The largest previously-unconstrained region is on the larger mass end of this plot, spanning several orders of magnitude both horizontally and vertically. However, there is also a somewhat difficult-to-see region above the ancient mica constraints which were previously unconstrained [33].
@@ II Energy transfer
First we need to estimate the effect of the collision between a dark matter bullet of mass mm and radius $r$ and a ‘regular’ baryonic spherical object with mass $M$ and radius $R$. While in general smaller asteroids and ring particles are not necessarily spherical, we will make this approximation for the sake of tractability—more complex geometries will require detailed numerical simulation. We will also assume that the dark matter is itself not destroyed or otherwise majorly affected by the impact.
When the dark matter collides with the object, it will elastically scatter with particles in a tube of radius $r$ as it passes through, until it is either stopped in the material or exits out the other side. As a simplified way to average over impact angles, we will take the tube length to be the average secant length in a sphere, $\frac{4}{3}R$, so that the volume of this tube is $V=\frac{4}{3}\pi Rr^{2}$. This also sets a limit on the maximum dark matter radius $r$ such that the dark matter still passes cleanly through the object:
@@
Enforcing this ensures the geometry of the collision is sufficiently simple, and for the scenarios we consider here $r>0.25R$ occurs in regions of parameter space which are already ruled out by other dark matter constraints and can thus be safely ignored. In principal there is no reason that we cannot look at $r>0.25R$ to extend our limits upwards, but more work would be required for the unusual geometry there.
The number of collisions $N$ the dark matter experiences as it passes through the object is then given by $N\simeq\rho_{o}V/m_{p}$, where $\rho_{o}$ is the density of the object and where we use the proton mass $m_{p}$ as the approximate mass of a nucleon. We will assume the dark matter is colliding with a reasonably dense macroscopic object such that $N\gg 1$. For simplicity let us assume that the particles in the object are stationary with respect to the dark matter and treat each collision as a linear elastic collision. The change in speed after one such collision is simply,
@@
Then if $X>1$, the dark matter would have lost a significant portion of its velocity transiting the object, and if $X\gg 1$, the dark matter can become captured within the object. The energy $E$ imparted to the object by this collision can then be estimated by taking the difference between the kinetic energy of the dark matter before and after the impact. If the dark matter is captured, this is simply the total kinetic energy before the impact. If the dark matter is not captured, we then estimate,
@@
We will say that when the dark matter is at a distance $x$ throughout the tube, any particles found within the length $y$ from the center of the dark matter sphere will be ejected into space at the initial dark matter velocity. This will only serve as an extremely rough estimate for the complicated physics of the real shock which would precede the dark matter. We assume that any material farther than $y$ will deposit their energy into the material of the object, contributing to its destruction. With some geometrical working, the volume of particles lost $V_{\rm lost}$ is given by,
$V_{\rm lost}=\int_{0}^{4/3R}\mathrm{d}x\pi y^{2}=\pi r^{3}\tan^{-1}(R/r)~,$ (5)
so that the fraction of energy lost by this ejecta is simply,
@@ III Catastrophic Destruction
The study of the catastrophic destruction of asteroids has a long and rich history [34, 35, 36, 37, 38, 39, 40]. This is not only motivated by a desire to better understand how to avoid Earth-impacting events [41, 42, 43], but also to understand the history and formation of our Solar System [44, 45, 46, 47, 48, 49]. The generic criteria for asteroid destruction is quantified with the ratio $Q^{*}$, the kinetic energy of the impacter divided by the target mass. For catastrophic asteroid destruction, $Q^{*}_{d}$ is defined as the threshold above which two conditions hold: 1), the largest remaining fragment should be smaller than half the mass of the original; and 2), less than half of these fragments should eventually re-accumulate. We will adopt this criteria for destruction.
There are two important regimes for asteroids. Below roughly 0.510.5-1km, asteroids are considered to be in the material strength regime, where the primary forces which must be overcome to destroy them relate to the material strength of the rock or ice. Above this, the asteroids are gravity dominated, and the energy required to overcome the gravitational binding energy is significantly larger than the material strengths.
However, destruction of gravity dominated asteroids is generally more difficult than merely overcoming the gravitational binding energy. Larger asteroids (from about 200200m to 1010km) are now thought to be predominantly rubble-pile asteroids [50], gravitationally bound collections of rocks and boulders of varying sizes and compositions. Evidence for the ubiquity of rubble-pile asteroids are based on observations of low asteroid spins above about one kilometer in radius [51, 52], direct measurements of densities and compositions of individual asteroids [53, 54, 55], the observed breakup of Shoemaker-Levy 9 over Jupiter [56, 57, 58], and the DART mission [59, 60, 61, 62, 63] which intentionally collided with the asteroid Dimorphos. Since rubble pile asteroids can be highly porous, the energy of the impacter is distributed inefficiently, resulting in larger $Q_{d}^{*}$ than a naive estimate which only requires the gravitational binding energy be overcome. In general, impacts on rubble-pile asteroids must be simulated numerically, and there is a high degree of variance in outcomes depending on the propagation of shocks through the specific geometry and composition of the rubble pile.
@@
S-type $10^{3}$ $10^{6}$ $10^{-3}$ $500$ 0.33-0.33 $1.65$
C-type $2\times 10^{3}$ $4\times 10^{5}$ $10^{-3}$ $500$ 0.25-0.25 $1.23$
@@
Although the density of asteroids is both size-dependent and has some variance amongst individuals, we used the benchmark values from Ref. [64]— for S-type asteroids we took the density to be $3.0$ g cm-3, and for C-types we used $1.7$ g cm-3. For the particles in Saturn’s rings we used the density of ice $0.92$ g cm-3, but with the best-fit $Q_{d}^{*}$ for S-type targets—while these are of course not designed for a pure ice target, earlier studies [35] show that the destruction threshold for ice and basalt do not differ by more than a factor of approximately two, with ice being the easier of the two to destroy.
The smallest observed asteroids are about 22m large, such as TC25 [65]—although this is not strictly an S- or C-type asteroid, we will use it as the benchmark smallest size. The largest S-type asteroids are Eunomia and Juno [64], around $270$ and 250250km respectively, whereas the largest C-type are Hygiea at around 430430km and Ceres at 940940km [64]. However, our constraints are not very sensitive to the largest mass objects, since the destruction criteria ceases to be met before we reach such massive asteroids.
The composition of Saturn’s rings [66, 67, 68] vary by subdivision, but generally the smallest observed particles are around a few mm and the largest on the order of 1010m. To the best of our knowledge, there is no current evidence for smaller particles than these in stable orbits [69, 70], but if there was, the constraints here could be improved. We use 11mm as the small-size cutoff for Saturn’s ring particles, based on the treatment of monomers in Ref. [71] which we will discuss in more detail in the following subsection.
Since $Q_{d}^{*}$ is generally smaller for S-type asteroids, they tend to provide marginally tighter constraints on the dark matter than the C-type asteroids. We thus ultimately use the S-type asteroid formula when deriving the constraints. Eq. 7 holds until the objects are smaller than about $1$ m, below which the curve takes a constant value. In Ref. [34] a cutoff on this estimate is imposed at 11cm, but in order to consider the smallest particles in Saturn’s rings, we will naively extend this to 11mm.
@@
$t^{-1}\simeq n_{DM}\sigma v_{DM}~,$ (10)
where we take the collision cross section to be $\sigma=\pi\left(R+r\right)^{2}\simeq\pi R^{2}$. For a given object of mass $M$, radius $R$, and lifetime $\tau$, if $t<\tau$ and the energy of the impact is sufficient to destroy it, we consider the dark matter constrained. These constraints are plotted in Fig 1.
@@ III.1 Ages of asteroids and ring particles
The observed ages of asteroids and planetary rings must then be carefully considered in order to place constraints. For larger asteroids, observations of color [72] consistently confirm ages in the billions of years, and indeed rubble-pile asteroids are very difficult to destroy by collisions with other asteroids [63]. For smaller asteroids, to the best of our knowledge, data on age is scarcer. We then must turn to simulations of the evolution of the asteroid belt, such as those in Ref. [47]. The smaller the asteroid, the easier they are to catastrophically destroy by collisions with other asteroids. By simulating the entire population, it is then possible to determine the average collisional lifetime of an asteroid of specific size, as given in Table 2. We interpolate this data to determine the lifetime $\tau$ which should be compared against the destruction timescale $t$.
@@
Table 2: Relation between radius of asteroid and collisional lifetime from Ref. [47].
Of course, this lifetime is just the average lifetime of an asteroid of a given size. It is presumably the case that some of the small asteroids observed today are significantly older than the ages given here, which would substantially improve our derived dark matter constraints. To show this, the final constraint plot includes the region that would be constrained if we did indeed find small asteroids which could be proven to have survived $1$ Gyr without major disruptions. This is plotted as a dashed boundary in Fig. 1.
@@ III.2 Conservativeness of estimate
$E_{\mathrm{vap.}}/E_{k}=\frac{2H}{\rho_{\mathrm{basalt}}v^{2}}\simeq 0.6~,$ (11)
where $\rho_{\mathrm{basalt}}=2.8$g cm-3 and the standard bullet test speed is v=5.5v=5.5km/s. The energy lost to vaporization (if completely vaporized) can therefore be a sizable fraction of the total kinetic energy of the impacter. In our scenario, in contrast, no energy would be lost to this vaporization.
Estimating the energy losses due to the geometry of the shocks is far more complicated, but we can get at least a rough sense of the estimation. One historically useful criteria [34, 35] for asteroid destruction used the fact that the velocity of the material shocks dropped with the distance squared in the material as the shock spread from the impact point. If the bullet impacted on one side of the material, one could then calculate how fast the shock is moving on the opposite side of the asteroid, at a distance $2R$, and if the velocity at that distant point was still larger than the escape velocity, then the asteroid could be considered disrupted. In our case, the energy is not just deposited at the impact site, but throughout the secant the dark matter bullet carves out. The most distant point from the average secant we considered here is perpendicular from its midpoint, at a distance approximately $1.75R$. Then the velocity of the shock at this point will be $1.3$ times the velocity of the shock at $2R$. If we instead considered a trajectory through the middle of the asteroid (rather than the average secant), the velocity would be four times larger. Combined with the previous estimate of the energy lost to vaporization, then, we could be underestimating the constraints by an order of magnitude or so.
@@ IV Orbit evaporation
If objects in our Solar System were moving through an isotropic background of dark matter, then there would be a consistent bias towards head-on collisions, regardless of the direction of motion of the object with respect to the Solar frame of reference. This would cause the objects to on average experience a drag force, provided the momentum transfer is sufficiently small in each collision. However, this situation does not occur for the objects in consideration here—the Solar System is moving through the galaxy at a rate approximately equal to the dark matter velocity, causing the dark matter background to be extremely anisotropic.
While this dark matter ‘wind’ could be ignored in the previous consideration of destruction (where only one collision is necessary to place constraints), it can no longer be ignored here. Not only is the most probable collision always occurring in the direction of the wind (rather than the direction of motion of the object itself), but the relative velocity in the headwind is significantly higher and so momentum transfer is larger. In the frame of reference of the galaxy, this will look like a drag force on objects in our Solar System, whereas in the Solar frame, it will look like the objects are being accelerated in the direction of the wind. It is important to remember that the orbital plane of the planets in our Solar System is not aligned with the axis of the Milky Way, being tilted around $60^{\circ}$ with respect to the Milky way plane. The dark matter wind is currently coming from the direction of the constellation Cygnus, at a declination close to $40^{\circ}$.
The Solar System orbits the galaxy with a period of approximately 230 million years. Thus on timescales of $\mathcal{O}(10~\rm Myr)$ or less we can treat the dark matter wind as a force acting on the object in a constant direction. Over longer timescales, the direction of the dark matter wind rotates through $2\pi$ as the Solar System orbits the Milky Way. If the dark matter collisions constituted a constant, continuous force on any object, this force would average out to zero over the course of one galactic orbit. As we will see, however, when collisions with the dark matter are infrequent enough to be discrete events (as is the case for macroscopic dark matter), the expected change in velocity after one galactic orbit is larger than zero. The intuition for this observation is to imagine that the collisions are changing the object’s velocity in a ‘random walk.’ In a random walk, the average final position remains at the origin, (over many trials), but the average displacement goes like the square root of the number of steps.
It is beyond the scope of this paper to treat the dark matter wind in full details, including the dark matter velocity distribution, careful integration over long timescales, and the resulting distribution of final velocities of the objects. We intend to treat this more thoroughly in forthcoming works. To make a simple order-of-magnitude estimate here, we will merely calculate the expected change in velocity after one full galactic orbit. We will then take that acceleration and use it as the average acceleration, even over non-integer numbers of orbits around the Milky Way. This will give us a rather conservative estimate of the final velocities of the objects, since over partial orbits of the galaxy, the directions of dark matter collisions are more aligned and so would in fact accelerate the object at a higher rate.
Ultimately we want to know how many collisions $k$ are required to increase the velocity by a factor $\delta$. For evaporation, escape velocity at a given radius is only a factor $\delta=\sqrt{2}$ larger than the speed of circular orbit, so $\delta=\sqrt{2}$ will be the key threshold. We will use this threshold as a conservative way to place constraints, since the evaporation of objects from their orbits would trivially contradict observations. However, it is certainly the case that we could place much more sensitive constraints than this—even a small disruption to the trajectories of Saturn’s rings, for instance, should have large observable consequences. We will leave this more sensitive computation to future work, noting that even our more crude constraints eliminate much of the meaningful parameter space available.
The number of collisions over some timespan $\tau$ is simply,
$k=\frac{\rho_{DM}}{m}\sigma v_{\rm rel}\tau~,$ (12)
where we again take the cross section $\sigma=\pi\left(R+r\right)^{2}\simeq\pi R^{2}$ and where $v_{\rm rel}$ is the relative velocity of the dark matter wind, which will assume for simplicity to be $v_{\rm rel}=2v_{\rm DM}$.
@@
$\vec{v}_{f}=\vec{v}_{o}+v_{rel}\frac{m}{M}\left(1-e^{-X}\right)\sum_{k}\hat{r}_{i}~,$ (13)
Where $\vec{v}_{o}$ is the object’s initial velocity. To account for the change in the dark matter wind, $\hat{r}_{i}$ are random unit vectors taken from the set of directions the dark matter wind points over one full galactic orbit (i.e., a two-dimensional circle). We are simplifying here by assuming that all collisions are happening in the direction of the dark matter wind, since it is by far the most probable case. A full calculation which averages over all collisions in three dimensions is beyond the scope of this simple estimation.
We will then make use of a handy result for the expected value of the magnitude of the sum of $k$ random, isotropic unit vectors [79]. When $k$ is large, for vectors in $\mathbb{R}^{d}$ the expected value of this magnitude approaches,
[…]
which evaluates to $\simeq 0.89\sqrt{k}$ in two dimensions, as we require here.
Before proceeding, it is important to point out the effect of this square root on the calculation. One might naively expect that the evaporation of Solar System bodies should be independent of dark matter mass, since a smaller mass implies less momentum transfer but proportionately more collisions. However, averaging over the isotropy of the dark matter and the resulting random walk of the object’s momentum is responsible for this $\sqrt{k}$ growth which ultimately breaks the degeneracy in dark matter mass. As the dark matter mass decreases, we easily see that $\vec{v}_{f}-\vec{v_{o}}\propto\sqrt{m}\rightarrow 0$ and we recover the required result that a continuous force rotated around $2\pi$ will not have impacted the final velocity of the object. Of course, it would be interesting to consider what the actual displacement of the object would look like over one full (or partial) orbit of the galaxy—this could lead to more sensitive constraints which we leave to a more detailed forthcoming analysis of this effect.
@@
$\sqrt{k}$ $\geq\frac{1}{0.89}\frac{M}{m}\frac{v_{o}}{v_{\rm rel}}\frac{\delta-1}{\left(1-e^{-X}\right)}~.$ (15)
$\simeq\frac{1}{0.89}\frac{v_{o}}{v_{\rm rel}}\frac{R^{2}}{r^{2}}\frac{\delta-1}{2}~.$ (16)
Where the second line takes $X<1$. Substituting Eq. 12 allows us to estimate a lower limit on the dark matter radius such that the orbit will be eventually evaporated:
$r^{2}\geq\frac{R}{0.89}\left(\frac{m}{\pi\rho_{DM}v_{\rm rel}\tau}\right)^{1/2}\frac{v_{o}}{v_{\rm rel}}\frac{\delta-1}{2}~.$ (17)
Now let us return to the regime when $X>1$. For simplicity let us assume that in this limit the dark matter is captured in the object and all of the momentum is transferred. Let us also assume that this a perfectly inelastic collision, with no losses from cratering or ejecta. The calculation follows similarly to the previous section, with the final velocity now being given by,
$\vec{v}_{f}=\frac{1}{M+km}\left(M\vec{v}_{o}+mv_{\rm rel}\sum_{k}\hat{r}_{i}~\right),$ (18)
@@
$\frac{\sqrt{k}\frac{m}{M}}{\delta\left(1+(.89^{2})k\frac{m}{M}\right)-1}\geq\frac{1}{0.89}\frac{v_{o}}{v_{\rm rel}}~,$ (19)
which does not have a beautiful solution for $k$. We can proceed by noting that for the objects we have in question, $(.89^{2})km<M$ is very easy to satisfy, since this limit can be simplified to,
$R>2.9\times 10^{-6}~\mathrm{cm}\left(\frac{\rho_{o}}{1\mathrm{gcm}^{-3}}\right)^{-1}\left(\frac{\tau}{1\mathrm{Gyr}}\right)~,$ (20)